# Distributed Coordination for Heterogeneous Non-Terrestrial Networks

### v2

@@ Abstract
To ensure global coverage and ubiquitous connectivity, the Non-terrestrial Network (NTN) technology has been regarded as a key enabling technology in the Sixth Generation (6G) network, which consists of the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellite. Even within the same tier, such as the space tier, the NTNs are developed based on different platforms, including Low Earth Orbit (LEO), Medium Earth Orbit (MEO), and Geostationary Earth Orbit (GEO). It is noted that the unique characteristics of various NTN platforms directly impact the design and implementation of NTNs, resulting in highly dynamic and heterogeneous networks. Therefore, distributed coordination among heterogeneous NTNs remains an important challenge. Although a distributed learning framework finds a wide range of applications by leveraging rich distributed data and computation resources, the explicit and systematic analysis of the individual layers’ challenges, and corresponding distributed coordination solutions in heterogeneous NTNs has not been proposed yet. In this paper, we first summarize the unique characteristics of each NTN platform and analyze the corresponding impact on its design and implementation. We then identify the communication challenges of heterogeneous NTNs in individual layers, where the potential delay-tolerant or delay-sensitive coordinated solutions are proposed accordingly. We further illustrate the multi-agent deep reinforcement learning (MADRL) algorithms tailored for the proposed coordinated solutions. Last but not least, we present a case study of the joint optimization problem on user scheduling and trajectory design in heterogeneous NTN containing tethered-UAV (T-UAV) and untethered-UAV (U-UAV), where the two-timescale multi-agent deep deterministic policy gradient (MADDPG) technique is developed to validate the effectiveness of distributed coordination.
Index Terms:
Distributed Coordination, Distributed Learning, Heterogeneous Network, Non-terrestrial Network
@@ I Introduction
With the improved coverage and resilience against terrestrial infrastructure failures, the Non-terrestrial Network (NTN) has been identified as one of the emerging usage scenarios in International Mobile Telecommunications 2030 (IMT-2030) for the Six Generation (6G) network, which has the potential to provide connections in post-disaster scenarios and remote areas [1]. Currently, the research in the NTN community focuses on three platforms including the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellites. It is noted that the unique characteristics of each platform (e.g., power supply and loading capability) play a critical role in the design and implementation of the NTN (e.g., transmission power and number of antennas) [2]. Therefore, the coordination of heterogeneous NTNs remains an important challenge to solve.
@@
    We further present emerging multi-agent deep reinforcement learning (MADRL) algorithms tailored for coordinated heterogeneous NTNs in Section IV, which are divided into delay-tolerant and delay-sensitive functions, respectively.
@@
    We present a case study of joint optimization on user scheduling and trajectory design in a heterogeneous NTN network composed of tethered-UAV (T-UAV) and untethered-UAV (U-UAV), where the multi-agent deep deterministic policy gradient (MADDPG) algorithm is utilized to validate the effectiveness of our proposed distributed coordination method. Finally, we conclude the paper in Section VI.
@@ II Unique Characteristics of Non-terrestrial Network Platforms
Fig. 1 shows the typical terrestrial 5G network consisting of a Radio Unit (RU), a Distributed Unit (DU), a Centralized Unit (CU), and the core network, where O-RAN CatA is adopted for the fronthaul split [6]. To achieve the tradeoff between fronthaul bandwidth and coordination, the beamforming procedure, inverse fast Fourier transform (iFFT), and cyclic prefix processing are integrated within the RU, where the user data symbols and precoding weights are transmitted over the fronthaul. In terms of the Size, Weight, and Power (SWaP), it is noted that the typical weights of macro massive MIMO RU, and Baseband Unit (BBU) are over 10 kg and 5 kg, respectively, where BBU consists of DU and CU. The average power consumption of RU and CU can reach 1.175 kWh and 0.325 kWh under a full workload, which can even reach 3.8 kWh with the three-sector setting.
@@
II-A Unmanned Aerial Vehicle
@@ II-A1 Untethered UAV
Taking DJI as an example, the flagship UAV platform Matrice 350 RTK achieves a loading capability of 2.7 kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above. Although DJI has customized a delivery UAV platform with a 40 kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users. Apart from that, the battery is originally designed to support the UAV flight with a typical capacity of 0.3 kWh (e.g., TB65 from DJI). Therefore the commercial untethered UAV can only be utilized to implement a single micro or pico network with RU, DU, and CU integrated as a complete payload.
@@ II-A2 Tethered UAV
To overcome the limitations above, the tethered UAV has been regarded as a promising solution, and it has been utilized for emergency communication in post-disaster areas. Based on the power supply provided by the tethered system, the tethered UAV achieves much stronger loading capability and endurance time, which is suitable to implement a single macro network, e.g., DG-M30 tethered UAV can carry 20 kg payloads hovering at 200 m for 24 hours. More importantly, the optical fiber provided by the tethered system enables flexible communication payload options, including RU, RU/DU, and RU/DU/CU. However, the tethered system limits the mobility of the UAV, which may lead to coverage degradation.
@@ III-A Physical Layer
More importantly, the uplink transmission performance should also be considered during cell selection, where the UE may experience poor uplink performance to certain cells due to limited transmission power. To solve the problem, a hierarchical coordination framework should be developed for cell selection, where the NTNs perform coordinated intelligent cell shaping using optimal SSB beams, and then each UE selects the cell based on its local characteristics including transmission power. One typical example is the intelligent cell shaping developed for terrestrial networks by Ericsson, where multiple BSs coordinate to optimize the cell coverage area based on reinforcement learning. It is noted that the cell selection procedure is delay-tolerant over the Telecom operator’s level.
Channel Estimation: Uplink channel estimation has been regarded as an important step in guiding downlink user data transmission, including scheduling in the MAC layer and precoding in the physical layer, where each UE is required to transmit the uplink pilot signal to the NTNs, i.e., Sounding Reference Signal (SRS). If the UE is associated with a higher tier NTN, higher transmission power is usually needed to compensate for the larger path loss. Therefore, in the vertical heterogeneous NTNs, UEs associated with lower-tier NTNs will experience serious pilot interference from the UEs associated with the higher-tier NTNs. Considering the much larger beam footprint in the higher-tier NTNs and limited beamforming capability at the UE, pilot interference will be more serious than that in traditional terrestrial networks.
@@
Power Allocation: Power allocation optimization has been regarded as an effective solution to maximize the channel capacity both in single-cell and multiple-cell scenarios. In the heterogeneous NTNs, cross-link interference (CLI) and remote interference (RI) emerge as two important challenges due to the LoS-dominated aerial channel, where CLI happens between adjacent cells with different TDD frame structures, and RI happens between very distant cells even with synchronized TDD frame structures. It is noted that the upper-tier NTNs usually adopt a much higher downlink transmission power than the lower-tier NTNs, which will cause strong CLI and RI to the lower-tier NTNs. To solve the CLI and RI in heterogeneous NTNs, especially for the downlink (DL) to uplink (UL) direction, a coordinated UL-DL power allocation scheme should be developed to enhance the UL signal strength and mitigate DL interference. Similar to precoding, the power allocation procedure is delay-sensitive and needs to be executed in each time slot.
@@ III-C Network Layer
Routing: Since diverse mobility patterns of satellites, HAPs, and UAVs lead to the dynamic backhaul topologies, it is difficult to maintain stable point-to-point LoS backhaul transmission between the NTN platforms and the core network on the ground. Therefore, the integrated access and backhaul (IAB) technology has been regarded as a promising solution for more robust backhaul transmission [11], which has been further extended to wireless access backhaul (WAB) at 3GPP Release 19. More importantly, the Backhaul Adaptation Protocol (BAP) layer is introduced in DU for routing, which makes it possible to achieve multi-hop backhaul transmission. However, the dynamic topology and heterogeneous communication capabilities make it challenging to find the optimal routing path. Therefore, the coordinated routing algorithm should be developed by considering both the local cell’s status and the neighboring cells’ status. Another challenge in the backhaul routing via IAB is the channel capacity of the aerial communication. Unlike traditional access links, the IAB link transmits cell-level traffic between two NTNs via an access link, which has a much higher capacity requirement. However, the propagation path is usually dominated by LoS in aerial links, which limits the application of spatial multiplexing to enhance channel capacity. One promising solution is the multi-connectivity technology to enable distributed intelligent backhaul traffic allocation on multiple routing paths. The routing table optimization is classified as a delay-tolerant function, which aims to keep the traffic relatively stable over a period of time and collect enough observations before making a new decision.
@@ IV-A Delay-tolerant Distributed Coordination
According to the O-RAN structure, the near real-time (near-RT) RAN intelligent controller (RIC) deployed in the CU can typically support the RAN control with latency from 10 ms to 1 s, and the non-RT RIC deployed in the core network can support the RAN control with latency over 1 s. Therefore, for delay-tolerant distributed coordination, the centralized server during training can be deployed in CU either onboard or on the ground, or even in the core network.
@@
  • Multi-hop IAB: Although the training of the routing optimization algorithm is delay-tolerant, the user data transmission may have stringent latency requirements. Therefore, multi-hop transmission of duplicate user data leads to increased transmission latency and channel capacity demands. To solve the problem, the hierarchical goal-oriented semantic communication can be integrated with the MADRL to reduce transmission overhead by prioritizing task-relevant information in each hop [14]. Additionally, the graph encoders, such as GNN, can also be incorporated into each agent’s observation, enhancing the adaptability of MADRL policies to topological changes.
@@ IV-B Delay-sensitive Distributed Coordination
It is noted that the cellular network has stringent numerology and frame structure, where most of the physical layer and MAC layer procedures need to be completed within every slot length. To satisfy the stringent latency requirement during coordination, the central server during training should be deployed in the DU, and a single DU should connect to multiple RUs for coordination. In this case, more functions in the physical layer and MAC layer are required to be moved to the RU.
  • Distributed MIMO: Since the heterogeneous NTNs are equipped with a distinct number of antennas, the observation space (e.g., dimensions of the estimated channel) and action (e.g., dimensions of the precoding matrix) space of different agents vary from each other, the distributed MADRL algorithm with heterogeneous agents including MADDPG should be investigated. Considering the channel sharing overhead, to fully enable DMIMO precoding without channel sharing, one possible solution is to perform channel charting for cell-edge UEs via multiple cells in advance. Then, each NTN reconstructs the complete channel for cell-edge UEs based on the local estimated channel and channel charting map.
@@
  • Coordinated Scheduling: To reduce the inter-cell interference in the MAC layer, each agent should learn to cooperatively schedule the UEs with minimal channel correlation on the same resource block. In this case, the reward in MADRL can be designed as the sum capacity in the scenario. However, to further support CJT scheme in DMIMO, the MADRL algorithm will incorporate the distributed scheduling consensus constraint, where the coordinated cells should schedule the cell-edge UEs on the same resource block for precoding.
@@ V Case Study
To solve this problem, we propose a two-timescale MADDPG (TTS-MADDPG) algorithm under the CTDE framework. Specifically, this algorithm involves two groups of agents for two tasks: the first group (all UAVs) is responsible for the short-timescale user scheduling task at each time slot, while the second group (all U-UAVs) handles the long-timescale trajectory control task at every 5 time slots due to UAV’s mobility constraints. During the offline centralized training, global state information is utilized by the central critic to assist learning, while each agent’s actor is trained using only local observations.
@@
The average downlink throughput results of our TTS-MADDPG algorithm during both training and testing phases are illustrated in Fig. 3, where we also include two benchmark schemes for the comparison: a round-robin (RR)-based scheduling method and a MADDPG-based scheduling approach. We conduct the simulation across 1000 episodes, with each episode consisting of 200 time slots, and the learning rates of actor and critic set to $10^{-4}$ and $10^{-3}$. In Fig. 3(a), for the converged performance, the TTS-MADDPG algorithm yields approximately 150% and 19% throughput gains over the RR-based and the MADDPG-based scheduling benchmarks, respectively, converging at around 175 Mbps. During testing, the TTS-MADDPG algorithm maintains stable performance that closely matches its converged performance during training, demonstrating strong generalization capability. In Fig. 3(b), with the TTS-MADDPG algorithm, T-UAV’s throughput declines slightly to 27.6 Mbps, as it learns to ensure sufficient scheduling on U-UAVs to serve the edge users and maximize the overall throughput. Moreover, each U-UAV achieves a noticeable throughput improvement to about 36 Mbps because of the optimal scheduling and real-time trajectory control.
@@
The above results demonstrate that the proposed MADRL-based framework enables effective distributed coordination among heterogeneous UAV-based NTN platforms. While the case study focuses on the UAV-based scenario, it also exhibits rich heterogeneity in terms of physical characteristics, communication patterns, and the operation timescales, which acts as a solid foundation for extension to broader NTN architectures involving HAPs and satellites and for tackling other complicated challenges. Moreover, in the extended NTN scenario involving a vast number of agents and multiple timescales, it is challenging to achieve sufficient coordination, further leading to slow and unstable convergence in the MADRL training. To address these issues, except for the coordinated solutions mentioned above, the MADRL can be further integrated with the grouped training and decentralized execution framework or the graph encoders to reduce computational load and enhance the convergence.
VI Conclusion

### v3

@@ Abstract
To achieve global coverage and ubiquitous connectivity, the non-terrestrial network (NTN) has been regarded as a key enabler in the sixth generation (6G) network, which includes uncrewed aerial vehicles (UAVs), high-altitude platforms (HAPs), and satellites. Since the unique characteristics of various NTN platforms strongly affect their implementation and lead to a highly dynamic and heterogeneous NTN scenario, achieving distributed coordination remains an important research direction. However, the explicit and systematic analysis of the individual layers’ challenges and corresponding distributed coordination solutions in heterogeneous NTNs has not been proposed yet. Therefore, in this paper, we summarize the unique characteristics of each NTN platform, identify communication challenges within individual layers, and propose potential delay-tolerant or delay-sensitive coordinated solutions accordingly. We further analyse the feasibility of leveraging multi-agent deep reinforcement learning (MADRL) algorithms to achieve the proposed coordinated solutions. Finally, we present a case study of the joint scheduling and trajectory optimization problem in heterogeneous NTN, where a two-timescale multi-agent deep deterministic policy gradient (TTS-MADDPG) algorithm is developed to validate the effectiveness of distributed coordination.
@@ I Introduction
With the improved coverage and resilience against terrestrial infrastructure failures, the non-terrestrial network (NTN) has been identified as one of the emerging usage scenarios in International Mobile Telecommunications 2030 (IMT-2030) for the Six Generation (6G) network, which has the potential to provide connections in post-disaster scenarios and remote areas [1]. Currently, the research in the NTN community focuses on three platforms, including the uncrewed aerial vehicle (UAV), high-altitude platform (HAP), and satellites. It is noted that the unique characteristics of each platform (e.g., power supply and loading capability) play a critical role in the design and implementation of the NTN (e.g., transmission power and number of antennas) [2]. Therefore, the coordination of heterogeneous NTNs remains an important challenge to solve.
@@
    We further present emerging multi-agent deep reinforcement learning (MADRL) algorithm frameworks tailored for coordinated heterogeneous NTNs in Section IV, which are divided into delay-tolerant and delay-sensitive solutions, respectively.
@@
    We present a case study of joint optimization on user scheduling and trajectory design in a heterogeneous NTN network composed of tethered-UAV (T-UAV) and untethered-UAV (U-UAV), where the two-timescale multi-agent deep deterministic policy gradient (TTS-MADDPG) algorithm is utilized to validate the effectiveness of our proposed distributed coordination method. Finally, we conclude the paper in Section VI.
@@ II Unique Characteristics of Non-terrestrial Network Platforms
Fig. 1 shows the typical terrestrial 5G network consisting of a radio unit (RU), a distributed unit (DU), a centralized unit (CU), and the core network, where open radio access network (O-RAN) CatA is adopted for the fronthaul split [6]. To achieve the tradeoff between fronthaul bandwidth and coordination, the beamforming procedure, inverse fast Fourier transform (iFFT), and cyclic prefix processing are integrated within the RU, where the user data symbols and precoding weights are transmitted over the fronthaul. In terms of the size, weight, and power (SWaP), it is noted that the typical weights of macro massive MIMO RU, and baseband unit (BBU) are over 10 kg and 5 kg, respectively, where BBU consists of DU and CU. The average power consumption of RU and CU can reach 1.175 kWh and 0.325 kWh under a full workload, which can even reach 3.8 kWh with the three-sector setting.
@@
II-A Uncrewed Aerial Vehicle
@@ II-A1 Untethered UAV
Taking DJI as an example, the flagship U-UAV platform Matrice 350 RTK achieves a loading capability of 2.7 kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above. Although DJI has customized a delivery UAV platform with a 40 kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users. Apart from that, the battery is originally designed to support the UAV flight with a typical capacity of 0.3 kWh (e.g., TB65 from DJI). Therefore the commercial untethered UAV can only be utilized to implement a single micro or pico network with RU, DU, and CU integrated as a complete payload.
@@ II-A2 Tethered UAV
To overcome the limitations above, the tethered UAV has been regarded as a promising solution, and it has been utilized for emergency communication in post-disaster areas. Based on the power supply provided by the tethered system, the tethered UAV achieves much stronger loading capability and endurance time, which is suitable to implement a single macro network, e.g., DG-M30 tethered UAV can carry 20 kg payloads hovering at 200 m for 24 hours. More importantly, the optical fiber provided by the tethered system enables flexible communication payload options, including RU, RU/DU, and RU/DU/CU. However, the tethered system limits the mobility of the T-UAV, which may lead to coverage degradation.
@@ III-A Physical Layer
More importantly, the uplink transmission performance should also be considered during cell selection, where the UE may experience poor uplink performance to certain cells due to limited transmission power. To solve the problem, a hierarchical coordination framework should be developed for cell selection, where the NTNs perform coordinated intelligent cell shaping using optimal SSB beams, and then each UE selects the cell based on its local characteristics including transmission power. One typical example is the intelligent cell shaping developed for terrestrial networks by Ericsson, where multiple BSs coordinate to optimize the cell coverage area based on reinforcement learning. It is noted that the cell selection procedure is delay-tolerant at the Telecom operator’s level.
Channel Estimation: Uplink channel estimation has been regarded as an important step in guiding downlink user data transmission, including scheduling in the MAC layer and precoding in the physical layer, where each UE is required to transmit the uplink pilot signal to the NTNs, i.e., sounding reference signal (SRS). If the UE is associated with a higher-tier NTN, higher transmission power is usually needed to compensate for the larger path loss. Therefore, in the vertical heterogeneous NTNs, UEs associated with lower-tier NTNs will experience serious pilot interference from the UEs associated with the higher-tier NTNs. Considering the much larger beam footprint in the higher-tier NTNs and limited beamforming capability at the UE, pilot interference will be more serious than that in traditional terrestrial networks.
@@
Power Allocation: Power allocation optimization has been regarded as an effective solution to maximize the channel capacity both in single-cell and multiple-cell scenarios. In the heterogeneous NTNs, cross-link interference (CLI) and remote interference (RI) emerge as two important challenges due to the LoS-dominated aerial channel, where CLI happens between adjacent cells with different time division duplex (TDD) frame structures, and RI happens between very distant cells even with synchronized TDD frame structures. It is noted that the upper-tier NTNs usually adopt a much higher downlink transmission power than the lower-tier NTNs, which will cause strong CLI and RI to the lower-tier NTNs. To solve the CLI and RI in heterogeneous NTNs, especially for the direction from downlink (DL) to uplink (UL), a coordinated UL-DL power allocation scheme should be developed to enhance the UL signal strength and mitigate DL interference. Similar to precoding, the power allocation procedure is delay-sensitive and needs to be executed in each time slot.
@@ III-C Network Layer
Routing: Since diverse mobility patterns of satellites, HAPs, and UAVs lead to dynamic backhaul topologies, it is difficult to maintain stable point-to-point LoS backhaul transmission between the NTN platforms and the core network on the ground. Therefore, the integrated access and backhaul (IAB) technology has been regarded as a promising solution for more robust backhaul transmission [11], which has been further extended to wireless access backhaul (WAB) at 3GPP Release 19. More importantly, the Backhaul Adaptation Protocol (BAP) layer is introduced in DU for routing, which makes it possible to achieve multi-hop backhaul transmission. However, the dynamic topology and heterogeneous communication capabilities make it challenging to find the optimal routing path. Therefore, the coordinated routing algorithm should be developed by considering both the local cell’s status and the neighboring cells’ status. Another challenge in the backhaul routing via IAB is the channel capacity of the aerial communication. Unlike traditional access links, the IAB link transmits cell-level traffic between two NTNs via an access link, which has a much higher capacity requirement. However, the propagation path is usually dominated by LoS in aerial links, which limits the application of spatial multiplexing to enhance channel capacity. One promising solution is the multi-connectivity technology to enable distributed intelligent backhaul traffic allocation on multiple routing paths. The routing table optimization is classified as a delay-tolerant function, which aims to keep the traffic relatively stable over a period of time and collect enough observations before making a new decision.
@@ IV-A Delay-tolerant Distributed Coordination
According to the O-RAN structure, the near real-time (near-RT) RAN intelligent controller (RIC) deployed in the CU can typically support the RAN control with latency from 10 ms to 1 s, and the non-RT RIC deployed in the core network can support the RAN control with latency over 1 s. Therefore, for delay-tolerant distributed coordination solutions, the centralized server during training can be deployed in CU either onboard or on the ground, or even in the core network.
@@
  • Multi-hop IAB: Although the training of the routing optimization algorithm is delay-tolerant, the user data transmission may have stringent latency requirements. Therefore, multi-hop transmission of duplicate user data leads to increased transmission latency and channel capacity demands. To solve the problem, the hierarchical goal-oriented semantic communication can be integrated with the MADRL to reduce transmission overhead by prioritizing task-relevant information in each hop [14] in this heterogeneous network. Additionally, the graph encoders, such as GNN, can also be incorporated into each agent’s observation, enhancing the adaptability of MADRL policies to topological changes.
@@ IV-B Delay-sensitive Distributed Coordination
It is noted that the cellular network has stringent numerology and frame structure, where most of the physical layer and MAC layer procedures need to be completed within every slot length. To satisfy the stringent latency requirement of delay-sensitive coordination solutions, the central server during training should be deployed in the DU, and a single DU should connect to multiple RUs for coordination. In this case, more functions in the physical layer and MAC layer are required to be moved to the RU.
  • Distributed MIMO: Since the heterogeneous NTNs are equipped with a distinct number of antennas, the observation space (e.g., dimensions of the estimated channel) and action (e.g., dimensions of the precoding matrix) space of different agents vary from each other, the distributed MADRL algorithm with heterogeneous agents, including MADDPG, should be investigated. Considering the channel sharing overhead, to fully enable DMIMO precoding without channel sharing, one possible solution is to perform channel charting for cell-edge UEs via multiple cells in advance. Then, each NTN reconstructs the complete channel for cell-edge UEs based on the local estimated channel and channel charting map.
@@
  • Coordinated Scheduling: To reduce the inter-cell interference in the MAC layer, each agent should learn to cooperatively schedule the UEs with minimal channel correlation on the same resource block. In this case, the reward in MADRL can be designed as the sum capacity in the scenario. However, to further support the CJT scheme in DMIMO, the MADRL algorithm will incorporate the distributed scheduling consensus constraint, where the coordinated cells should schedule the cell-edge UEs on the same resource block for precoding.
@@ V Case Study
To solve this problem, we propose the TTS-MADDPG algorithm under the CTDE framework. Specifically, this algorithm involves two groups of agents for two tasks: the first group (all UAVs) is responsible for the short-timescale user scheduling task at each time slot, while the second group (all U-UAVs) handles the long-timescale trajectory control task at every 5 time slots due to UAV’s mobility constraints. During the offline centralized training, global state information is utilized by the central critic to assist learning, while each agent’s actor is trained using only local observations.
@@
The average downlink throughput results of our TTS-MADDPG algorithm during both training and testing phases are illustrated in Fig. 3, where we also include two benchmark schemes for comparison: a round-robin (RR)-based scheduling method and a MADDPG-based scheduling approach. We conduct the simulation across 1000 episodes, with each episode consisting of 200 time slots, and the learning rates of actor and critic set to $10^{-4}$ and $10^{-3}$. In Fig. 3(a), for the converged performance, the TTS-MADDPG algorithm yields approximately 150% and 19% throughput gains over the RR-based and the MADDPG-based scheduling benchmarks, respectively, converging at around 175 Mbps. During testing, the TTS-MADDPG algorithm maintains stable performance that closely matches its converged performance during training, demonstrating strong generalization capability. In Fig. 3(b), with the TTS-MADDPG algorithm, T-UAV’s throughput declines slightly to 27.6 Mbps, as it learns to ensure sufficient scheduling on U-UAVs to serve the edge users and maximize the overall throughput. Moreover, each U-UAV achieves a noticeable throughput improvement of about 36 Mbps because of the optimal scheduling and real-time trajectory control.
@@
The above results demonstrate that the proposed MADRL-based framework enables effective distributed coordination among heterogeneous UAV-based NTN platforms. While the case study focuses on the UAV-based scenario, it also exhibits rich heterogeneity in terms of physical characteristics, communication patterns, and the operation timescales, which acts as a solid foundation for extension to broader NTN architectures involving HAPs and satellites and for tackling other complicated challenges.
VI Conclusions and Future Work